OpenStudy (anonymous):

PLEASE HELP D: I really need help with about 8 questions????

OpenStudy (anonymous):

I can try and help with some but I don't know...

OpenStudy (anonymous):

yay! Okay so I'll give you the first 3.......

OpenStudy (anonymous):

Ok for the first question here: Answer: (A) -> [ 98h^3 ] ^(1/2) -> [ (49h^2) x (2h)]^ (1/2) the square root of 49 =7 and the square root of h^2 = h -> 7h x [ (2h) ]^(1/2)

OpenStudy (anonymous):

THANK YOU :D

OpenStudy (anonymous):

Can you answer any more?

OpenStudy (anonymous):

Yes...I'm working on it now

OpenStudy (anonymous):

Yay!

OpenStudy (ivettef365):

on 11 just combine like terms you can add (3-9)sqrt11 and sqrt6 stays as it is

OpenStudy (anonymous):

The answer to second question: B. Just combine like terms and I did this on my calculator too

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

That's all I can do for now...but I'll tag some people...maybe they can help you, cause' I got to go @BulletWithButterflyWings @chaylacryder @Kitt020912 @kyrsten17 @uri there ya go Bye!

OpenStudy (anonymous):

okay thanks :)

OpenStudy (anonymous):

Those are the questions :)

OpenStudy (anonymous):

I got 10.11 answered

OpenStudy (anonymous):

All I need now is 12,13,15,16, and 20 :)

OpenStudy (phi):

for question 12, did you try factoring 52 and 117 ?

OpenStudy (anonymous):

Do you mean 13?

OpenStudy (phi):

yes, I mean Q13.

OpenStudy (anonymous):

what is factoring again?

OpenStudy (phi):

breaking up a number into 2 numbers that multiply together. Hint: one of the numbers will be 13 (if this will simplify)

OpenStudy (anonymous):

ohh okay!

OpenStudy (anonymous):

so 4*13 and 9*13?

OpenStudy (phi):

yes. now use the idea \[ \sqrt{4 \cdot 13} = \sqrt{4} \sqrt{13} \]

OpenStudy (anonymous):

is √13√4 the answer?

OpenStudy (phi):

it's on the way. Are you saying you don't know the square root of 4 ?

OpenStudy (anonymous):

its 2.

OpenStudy (phi):

so sqrt(4) sqrt(13) is the same as 2 sqrt(13) what about \[ \sqrt{9} \sqrt{13} \]

OpenStudy (anonymous):

3 sqrt (13)

OpenStudy (phi):

so your problem can be written as \[ \sqrt{52} - \sqrt{13} + \sqrt{117} \\ 2\sqrt{13} - 1 \sqrt{13} + 3 \sqrt{13} \] if you have 2 square roots of 13, take away 1, and add 3, how many square roots of 13 do you have ?

OpenStudy (anonymous):

4!

OpenStudy (phi):

so the answer is \[ 4 \sqrt{13} \]

OpenStudy (anonymous):

Yay! Thanks!

OpenStudy (phi):

for Q15, you can change the order of the multiplies \[ 2 \sqrt{15 k} \cdot 6k\sqrt{3} \\ 2 \cdot 6\cdot k \cdot \sqrt{15k} \sqrt{3} \] you should simplify 2*6*k you should factor the 15.

OpenStudy (anonymous):

1*3*k. Get rid of the 1. 15= 5*3

OpenStudy (phi):

2*6*k is not 1*3*k write down what you have so far

OpenStudy (anonymous):

so 12*k?

OpenStudy (phi):

yes, or just 12k (in algebra people leave out the multiply sign when you have a number and a letter.... of course this won't work for two numbers because 2*2 written as 22 is wrong) but write down what you have, including the radicals.

OpenStudy (anonymous):

Okay I am just going to guess on the rest of these questions... Thanks for the help though!! :)

OpenStudy (phi):

It is one thing to not know this... it is another to be lazy

OpenStudy (anonymous):

I know, I just have too much to do and I started late.